OpenStudy (anonymous):

Could someone please check me and tell me if I'm right?! A rectangle's length is 4 feet more than its width. If the area of the rectangle is 396 square feet, what is its width, in feet? I got 18, could someone tell me if I'm correct, and if not, tell me what the answer is or help guide me to it?

OpenStudy (anonymous):

it looks right to me

OpenStudy (ybarrap):

L = 4 + W A = L*W=368 = (4+W)*W W^2 + 4W -368 =0 W= 2 (sqrt(93)-1) = 17.28, using only the positive root- http://www.wolframalpha.com/input/?i=W%5E2+%2B+4W+-368

OpenStudy (ybarrap):

http://www.wolframalpha.com/input/?i=W^2+%2B+4W+-396 Ans. 18 as you concluded.